# Parallel Heuristic Exploration for Additive Complexity Reduction in Fast Matrix Multiplication

### v1

@@ 4.2.2 Equal Performance
Table 2 shows representative schemes where the proposed method matches GP-40. Both methods achieve substantial reductions from naive counts.
Table 2: Schemes with equal addition counts (proposed method matches GP-40)
@@
$m\times n\times p$ 5 40 $m\times n\times p$ 5 40
$2\times 2\times 2$ 7 24 15 15 15 15 $2\times 4\times 10$ 65 340 223 222 222 222
$2\times 2\times 3$ 11 20 20 20 20 20 $2\times 5\times 10$ 80 416 275 273 273 273
$2\times 2\times 4$ 14 36 31 31 31 31 $2\times 6\times 6$ 57 326 231 228 228 228
$2\times 2\times 5$ 18 38 33 33 33 33 $2\times 7\times 7$ 77 452 323 320 320 320
$2\times 2\times 6$ 21 54 44 44 44 44 $2\times 8\times 8$ 100 608 427 424 424 424
$2\times 2\times 7$ 25 56 46 46 46 46 $3\times 3\times 3$ 23 97 60 60 60 60
$2\times 2\times 8$ 28 72 57 57 57 57 $3\times 3\times 4$ 29 134 93 92 92 92
$2\times 2\times 9$ 32 74 59 59 59 59 $3\times 3\times 5$ 36 193 125 123 123 123
$2\times 2\times 10$ 35 90 70 70 70 70 $3\times 3\times 7$ 51 279 155 154 154 154
$2\times 2\times 11$ 39 92 72 72 72 72 $3\times 3\times 8$ 58 275 170 169 169 169
$2\times 2\times 12$ 42 108 83 83 83 83 $3\times 3\times 10$ 72 386 217 214 214 214
$2\times 2\times 13$ 46 110 85 85 85 85 $3\times 3\times 11$ 79 493 296 287 286 286
$2\times 2\times 14$ 49 126 96 96 96 96 $3\times 3\times 13$ 94 593 341 335 334 334
$2\times 2\times 15$ 53 128 98 98 98 98 $3\times 3\times 15$ 108 579 309 305 305 305
$2\times 2\times 16$ 56 144 109 109 109 109 $3\times 3\times 16$ 115 862 424 410 407 407
$2\times 3\times 3$ 15 58 46 46 46 46 $3\times 4\times 4$ 38 194 136 133 133 133
$2\times 3\times 4$ 20 82 58 58 58 58 $3\times 4\times 5$ 47 277 170 161 161 161
$2\times 3\times 5$ 25 110 72 71 71 71 $3\times 4\times 7$ 64 446 252 251 249 249
$2\times 3\times 6$ 30 116 81 81 81 81 $3\times 4\times 8$ 74 461 272 267 267 267
$2\times 3\times 7$ 35 140 100 99 99 99 $3\times 4\times 9$ 84 535 324 317 316 316
$2\times 3\times 8$ 40 164 104 104 104 104 $3\times 4\times 11$ 102 641 387 384 381 381
$2\times 3\times 9$ 45 174 116 116 116 116 $3\times 4\times 12$ 111 721 424 412 411 411
$2\times 3\times 11$ 55 222 146 145 145 145 $3\times 4\times 13$ 121 804 471 456 455 455
$2\times 3\times 12$ 60 232 151 151 151 151 $3\times 4\times 14$ 128 896 466 460 458 458
$2\times 3\times 14$ 70 280 181 180 180 180 $3\times 4\times 16$ 148 1043 532 518 518 518
$2\times 3\times 15$ 75 307 192 192 192 192 $3\times 5\times 8$ 94 584 295 295 295 295
$2\times 3\times 16$ 80 328 196 196 196 196 $3\times 7\times 8$ 128 930 462 450 450 450
$2\times 4\times 4$ 26 130 93 92 92 92 $3\times 8\times 8$ 148 1020 519 511 511 511
$2\times 4\times 5$ 33 184 114 112 112 112 $4\times 5\times 5$ 76 530 315 300 299 299
$2\times 4\times 6$ 39 202 138 136 136 136
These 59 schemes represent cases where both methods find the same (likely optimal or near-optimal) solution. In most of these cases, GP-5 and GP-40 produce identical addition counts, suggesting these values may indeed be optimal for the given algebraic form. The consistent 35-45% reduction from naive counts demonstrates the effectiveness of CSE optimization.
@@ 4.2.3 Inferior Performance
Table 3 shows the 2 schemes where GP-40 achieves lower addition counts than the proposed method. For all of these 2 schemes the proposed method still beats the default GP-5 configuration.
@@
The performance gap grows with scheme size. The largest differences are 43 additions for $(6,6,10:252)$. However, GP-40 requires running the optimization with 40 different $\alpha$ parameter values, which for large schemes can take tens of hours of computation time. In contrast, the parallel heuristic CPU search completes in 3-4 hours.
These results highlight the combinatorial difficulty of additive complexity reduction. While exhaustive parameter sweeps (GP-40) can find better solutions for very large schemes, they come at significant computational cost. To match GP-40’s results, the proposed method would likely need substantially more computation time. The current implementation provides a favorable trade-off, achieving near-optimal results much faster for most schemes, and still outperforming the practical default configuration (GP-5) in most challenging cases.
@@ 4.2.4 Component-Wise Analysis
$m\times n\times p$ Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min
@@
This analysis suggests potential for a hybrid approach: combine the best-optimized U, V and W components from different optimization runs to achieve even lower total addition counts. For example, taking the minimum from each column in Table 4 would reduce totals below what either method achieves alone.
@@ 4.2.5 Summary Statistics
Across the 164 ternary-coefficient schemes tested, the proposed parallel method demonstrates consistent improvement over the state-of-the-art:
  • 103 schemes (62.8%): proposed method achieves lower addition counts;
  • 59 schemes (36.0%): proposed method matches GP-40 results;
  • 2 schemes (1.2%): GP-40 achieves better addition counts.
For most cases, the parallel heuristic search provides equal or better results than the exhaustive GP-40 method while being significantly faster. On moderate-sized schemes, the proposed approach typically completes in minutes, while GP-40 can take hours.
For the largest schemes (typically rank >250>250), the pattern changes. The proposed method may require several hours and still yield slightly higher addition counts than GP-40. This highlights the combinatorial difficulty of additive reduction: random exploration works well when the search space is manageable but becomes less effective as dimensionality grows.
Despite this limitation, the proposed method consistently outperforms the practical default configuration GP-5, achieving lower addition counts on 115 schemes (70.1%). This makes it a valuable tool for rapid exploration and optimization of matrix multiplication schemes, especially during algorithm development where fast iteration is important.
@@ 5.1 Limitations
  • Combinatorial complexity: as scheme size grows, the search space expands exponentially. Although the parallel tool explores thousands of paths, it cannot guarantee optimality. Results show diminishing returns on very large schemes, where exhaustive methods like GP-40 keep an advantage.
  • Heuristic dependence: the quality of results relies on the set of heuristics used. While the current heuristics work well for smaller schemes, larger schemes might benefit from different heuristic designs.
  • Ternary coefficient limitation: the method is developed and evaluated only for schemes with coefficients in $\{-1,0,1\}$. Its effectiveness on schemes with other coefficient sets (such as integer $\mathbb{Z}$ or rational $\mathbb{Q}$) remains untested.
@@ 6 Conclusion
This paper presents a parallel heuristic tool for reducing additive complexity in fast matrix multiplication. The method addresses the combinatorial challenge of common subexpression elimination by running many optimization processes at once, each using different selection strategies.
The main contributions are:
  • The Greedy-Intersections heuristic, which estimates the value of substitutions more quickly without expensive trial evaluations.
  • A parallel tool in both GPU (CUDA) and CPU (OpenMP) versions that explores the search space using diverse heuristics and shares promising partial solutions.
  • Evaluation on 164 ternary-coefficient schemes showing lower addition counts on 103 schemes, matching on 59, and worse on 2 compared to the state-of-the-art Greedy-Potential method.
The results show that parallel heuristic search, using randomness and multiple strategies, can find better reductions than purely greedy methods while being much faster. The tool offers a good balance between quality and speed, making it useful for algorithm development.
Analysis of $U$, $V$ and $W$ components separately reveals an important insight: different optimization approaches can find better reductions in different parts of the same scheme. Neither method consistently beats the other across all three components, suggesting potential for hybrid approaches that combine the best-optimized components from multiple runs.
For the largest schemes, the problem remains difficult, and exhaustive methods keep a small advantage at higher cost. Still, the proposed tool beats the practical default (GP-5) on 115 out of 164 schemes.
All software and optimized schemes are open source, providing a resource for further work on automating addition reduction in matrix multiplication algorithms. The component-wise analysis data in particular may guide future hybrid optimization strategies.

### v2

@@ 4.2.3 Equal Performance
The proposed method achieves the same addition count as GP-40 for 45 schemes. These results are summarized in two categories, based on the rank of the scheme.
Table 3 shows cases where the scheme’s rank matches the current best known in the literature. For many of these schemes, no previously published addition count was found to serve as a direct benchmark, apart from those cataloged by Martensson et al. [maartensson2025number]. Therefore, when the proposed method matches the result from the state-of-the-art GP-40 tool on these schemes, it confirms that both approaches converge to the same high-quality solution. This shared result likely represents the current best-known (and possibly optimal) addition count for that particular matrix multiplication algorithm.
Table 3: Equal addition counts (proposed method matches GP-40) for schemes with best-known rank
Format Rank Best Naive GV GP-5 GP-40 Proposed
$m\times n\times p$ $r$ known
$2\times 3\times 3$ 15 48 [maartensson2025number] 58 46 46 46 46
$2\times 3\times 4$ 20 99 [maartensson2025number] 82 58 58 58 58
$2\times 3\times 5$ 25 103 [maartensson2025number] 110 72 71 71 71
$2\times 3\times 6$ 30 - 116 81 81 81 81
$2\times 3\times 7$ 35 - 140 100 99 99 99
$2\times 3\times 8$ 40 - 164 104 104 104 104
$2\times 3\times 9$ 45 - 174 116 116 116 116
$2\times 3\times 11$ 55 - 222 146 145 145 145
$2\times 3\times 12$ 60 - 232 151 151 151 151
$2\times 3\times 14$ 70 - 280 181 180 180 180
$2\times 3\times 15$ 75 - 307 192 192 192 192
$2\times 3\times 16$ 80 - 328 196 196 196 196
$2\times 4\times 4$ 26 173 [maartensson2025number] 130 93 92 92 92
$2\times 4\times 6$ 39 - 202 138 136 136 136
$2\times 8\times 8$ 100 - 608 427 424 424 424
$3\times 3\times 3$ 23 60 [stapleton202560] 97 60 60 60 60
$3\times 3\times 4$ 29 105 [maartensson2025number] 134 93 92 92 92
$3\times 3\times 5$ 36 176 [maartensson2025number] 193 125 123 123 123
$3\times 4\times 4$ 38 198 [maartensson2025number] 194 136 133 133 133
$3\times 4\times 5$ 47 276 [maartensson2025number] 277 170 161 161 161
$3\times 5\times 10$ 115 - 730 391 380 380 380
$4\times 5\times 5$ 76 451 [maartensson2025number] 530 315 300 299 299
Table 4 presents schemes where the rank is above the best known. For these, the addition count is optimized for a non-optimal multiplication count.
Table 4: Equal addition counts (proposed method matches GP-40) for schemes with non-optimal rank
@@
$m\times n\times p$ $r$ 5 40 $m\times n\times p$ $r$ 5 40
$2\times 4\times 5$ 33 184 114 112 112 112 $3\times 4\times 7$ 64 446 252 251 249 249
$2\times 4\times 10$ 65 340 223 222 222 222 $3\times 4\times 8$ 74 461 272 267 267 267
$2\times 5\times 10$ 80 416 275 273 273 273 $3\times 4\times 9$ 84 535 324 317 316 316
$2\times 6\times 6$ 57 326 231 228 228 228 $3\times 4\times 11$ 102 641 387 384 381 381
$2\times 7\times 7$ 77 452 323 320 320 320 $3\times 4\times 12$ 111 721 424 412 411 411
$3\times 3\times 7$ 51 279 155 154 154 154 $3\times 4\times 13$ 121 804 471 456 455 455
$3\times 3\times 8$ 58 275 170 169 169 169 $3\times 4\times 14$ 128 896 466 460 458 458
$3\times 3\times 10$ 72 386 217 214 214 214 $3\times 4\times 16$ 148 1043 532 518 518 518
$3\times 3\times 11$ 79 493 296 287 286 286 $3\times 5\times 8$ 94 584 295 295 295 295
$3\times 3\times 13$ 94 593 341 335 334 334 $3\times 7\times 8$ 128 930 462 450 450 450
$3\times 3\times 15$ 108 579 309 305 305 305 $3\times 8\times 8$ 148 1020 519 511 511 511
$3\times 3\times 16$ 115 862 424 410 407 407
In most of these 45 cases, GP-5 and GP-40 produce identical addition counts, suggesting these values may represent optimal or near-optimal solutions for the given algebraic form. The consistent 35–45% reduction from naive counts across these schemes further demonstrates the effectiveness of CSE optimization.
@@ 4.2.4 Inferior Performance
For only 2 of the 149 tested schemes does the exhaustive GP-40 method find a lower addition count than the proposed parallel heuristic. These cases are shown in Table 5. Notably, for both schemes, the proposed method still achieves a lower count than the practical default GP-5 configuration.
@@
The performance gap is most pronounced for very large schemes. The largest difference is 43 additions for $(6,6,10:252)$. Achieving this result with GP-40 requires evaluating 40 different $\alpha$ parameter values – a process that can take tens of hours for such large schemes. In contrast, the parallel heuristic CPU search completes in 3–4 hours.
These results highlight the persistent combinatorial difficulty of additive complexity reduction. While exhaustive search can find slightly better solutions for the very largest schemes, it is much slower. The proposed method offers a better trade-off: it finds near-optimal solutions much faster for almost all schemes. It achieves near-optimal results significantly faster for the vast majority of schemes and still outperforms the practical default configuration (GP-5) in most challenging cases.
@@ 4.2.5 Component-Wise Analysis
$m\times n\times p$ $r$ Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min
@@
This analysis suggests potential for a hybrid approach: combining the best-optimized $U$, $V$ and $W$ components from different optimization runs could achieve even lower total addition counts than any single method. Notably, for the schemes $(4,7,8:164)$ and $(4,8,8:182)$ – which have optimal ranks – this component-wise minimum creates new SOTA totals of 689 and 793 additions, respectively. These improve upon the totals found by either method alone for these schemes.
@@ 4.2.6 Summary Statistics
Across the 149 ternary-coefficient schemes tested, the proposed parallel method shows clear improvements over the current best method:
  • 102 schemes (68.4%): proposed method achieves lower addition counts (including 57 new SOTA for optimal-rank schemes).
  • 45 schemes (30.2%): proposed method matches GP-40 results.
  • 2 schemes: GP-40 achieves better addition counts.
For the majority of cases, the parallel heuristic search provides equal or better results than the exhaustive GP-40 method while being significantly faster. On moderate-sized schemes, the proposed approach typically completes in minutes, while GP-40 can take hours.
For the very largest schemes (rank >250>250), the pattern changes. The proposed method may require several hours and can yield slightly higher addition counts than GP-40. This highlights the combinatorial difficulty of additive reduction: heuristic exploration works well when the search space is tractable but faces greater challenges as dimensionality grows.
Despite this limitation on the most extreme cases, the proposed method consistently outperforms the practical default configuration GP-5, achieving lower addition counts on 114 schemes (76.5%). This performance, combined with its ability to establish new SOTA records, makes it a valuable tool for rapid exploration and optimization of matrix multiplication schemes.
@@ 5.1 Limitations
  • Combinatorial complexity: as scheme size grows, the search space expands exponentially. Although the parallel tool explores many paths, it cannot guarantee optimality. Results show diminishing returns on large schemes, where exhaustive methods like GP-40 retain an advantage.
  • Heuristic dependence: the quality of results relies on the set of heuristics used. While the current heuristics work well for smaller schemes, larger schemes with more complex expression graphs might benefit from different heuristic designs or adaptive strategies that learn during optimization.
  • Coefficient set and scope: the method is developed and evaluated only for schemes with ternary coefficients in $\{-1,0,1\}$. Its effectiveness on schemes with other coefficient sets (such as integer $\mathbb{Z}$ or rational $\mathbb{Q}$) remains untested. Furthermore, this work focuses exclusively on minimizing the arithmetic operation count. Practical implementation performance also depends on factors such as memory access patterns and cache utilization, which are beyond the scope of this current optimization.
@@ 5.3 Future Work
  • Practical implementation and analysis: future work includes evaluating the practical performance of the optimized schemes. This involves analyzing their memory access patterns, conducting numerical stability experiments, and measuring actual runtime on target hardware to validate the real-world benefit of the reduced addition counts.
@@ 6 Conclusion
This paper introduced a parallel heuristic approach for reducing the additive complexity of fast matrix multiplication algorithms with ternary coefficients ($\{-1,0,1\}$). To overcome the combinatorial nature of common subexpression elimination, the method executes many independent optimization processes, guided by a diverse set of strategies including the novel Greedy-Intersections heuristic.
Evaluated on 149 known ternary-coefficient schemes, the approach demonstrates that massive parallel exploration is highly effective. It achieves lower addition counts than the exhaustive state-of-the-art method (GP-40) on 102 schemes (68.4%), matches it on 45 (30.2%), and is outperformed on only 2. Notably, this includes establishing 57 new state-of-the-art addition counts for schemes with optimal rank, directly improving the total arithmetic cost of the best-known algorithms for these matrix dimensions.
The analysis revealed that different optimization methods often excel on different parts ($U$, $V$ and $W$) of a scheme, pointing to a clear opportunity for future hybrid optimizers. While exhaustive search retains an advantage for the very largest schemes, the parallel heuristic provides a superior balance of speed and solution quality for the vast majority of cases, making it a practical tool for algorithm exploration and development.
All software, results, and optimized schemes are released as open source, providing a foundation for further research and a valuable resource for the computational algebra community.